I can't figure out these questions.
It is approximately 480 miles from Los Angeles, CA, to San Francisco, CA. Allowing for various traffic conditions, a driver can average approximately 60 miles per hour.
Write a linear function that expresses the distance treveled, d, as a function of time, t.

How far have you traveled after 4 hours?
Write a linear function that expresses the distance to b e traveled to reach San Francisco, s as a function of time, t.

Answer by Nate(3500)

Write a linear function that expresses the distance treveled, d, as a function of time, t.
t=d/60
How far have you traveled after 4 hours?
d=60t
d=60(4)
d=240 miles
Write a linear function that expresses the distance to be traveled to reach San Francisco, s as a function of time, t.
t=480-60x where x is the hours you already traveled